# EARN Fairness: Explaining, Asking, Reviewing and Negotiating Artificial Intelligence Fairness Metrics Among Stakeholders

### v1

@@ 6. Discussion
In this section, we will first discuss the limitations of our work and future research. Further, we will discuss the implications of our findings in the credit rating/loan scenario and the generalization of applying the EARN framework to other scenarios.
@@ 6.1. Limitations and Future Work
6.2. Fairness in the Credit Rating Scenario
Our work focused on one application domain in financial services and thus allowed us to advance our understanding of how decision subjects view fairness in this realm.
Whether based on weighted rankings of metric preferences or the frequency of Top-1 choices, stakeholders’ top 3 preferences were Conditional Statistical Parity, Counterfactual Fairness, and Consistency. This echoes findings from Nakao et al.’s study on loan scenarios (Nakao et al., 2023), where they observed loan officers tended to prioritize individual fairness. Similarly, Saxena et al. (Saxena et al., 2019) surveyed ordinary people in loan scenarios and also revealed that fairness metrics must account for the similarity among individuals, and thus Counterfactual Fairness and Consistency should be included.
Our results also emphasized the importance of considering non-protected features (e.g., employment) in AI fairness for credit rating. In this domain, employment status may, in the eyes of stakeholders, unfairly affect individuals from lower socioeconomic backgrounds. However, these must be carefully balanced with legally protected features (e.g., gender), to avoid discrimination. It could be the case that stakeholders focused on Conditional Statistical Parity because it served as an ideal balance between important non-protected features and protected features.
In this domain, stakeholders preferred subgroup fairness over simple group fairness. This means considering multiple legal requirements for protected attributes (e.g., gender and age) and, given the stakeholder views we gathered, also integrating non-protected features. However, this will bring with it issues with data scarcity in these groups, which will make comparisons difficult.
Interestingly, when teams negotiated the preferred fairness metric, we observed a common pattern where all teams tended to merge diverse preferences to form a more comprehensive and inclusive definition of AI fairness to resolve differences, despite using different negotiation strategies. That is, all teams agreed on a a dual focus on the absence of prejudice or favoritism toward both individuals and groups in the credit rating scenario. Our findings underscore the importance of integrating varied fairness considerations in this domain.
6.3. Generalizing the EARN Fairness Framework
We argue that tackling AI fairness transcends the bounds of solely technical, AI expert-driven approaches, evolving into a broader, human-centered concern. Our work demonstrates the feasibility of our approach. We will discuss how the EARN Fairness framework could be generalized to diverse contexts and scaled up to larger stakeholder groups.
@@ 6.3.1. Explaining Fairness Metrics
As mentioned in Section 2, most research indirectly explores metrics that resonate with users’ fairness perceptions (Srivastava et al., 2019; Nakao et al., 2022; Taka et al., 2023; Saxena et al., 2019). Instead, our framework shares a similar philosophy with the work of Cheng et al. (Cheng et al., 2021), using the FEE interactive system to reveal the underlying logic of metrics, and thus enabling users to decide which fairness metric aligns best with their expectations. Although our FEE tool only covered a subsection of metrics, this could be easily extended to cover more and different fairness metrics, for a range of features extending beyond protected and legitimate features. Our tool also allows researchers to extend it to different application scenarios by changing datasets or incorporating other AI models’ fairness outcomes.
@@
6.3.2. Eliciting Personal Metric Preferences and Mitigating Fairness Issues
@@
From the perspective of personal preferences, we found that stakeholders without AI expertise preferred fairness metrics significantly different from those commonly used by AI experts in industry and academia, such as demographic parity, equal opportunity, and equalized odds (Xu, 2022; Bajracharya et al., 2023). Similarly, Nakao et al. (Nakao et al., 2023) found that loan officers’ preferences diverged from data scientists’. This highlights the need to involve a broader base of stakeholders in fairness metric decisions. However, even within the same type of stakeholders, in this case, decision subjects, we observed diverse understandings of fairness, not only in metrics but also in fairness categories. This finding echoes previous research (Lee et al., 2020; Landers and Behrend, 2023; Cheng et al., 2021).
This leads to the question of what to do with personal preferences once we have elicited them, in order to mitigate fairness issues. One approach might be to incorporate some simple voting for metrics and features, similar to the weighting that we have used, with the ”winning” metric and features being the ones implemented (Zhang et al., 2023). We might also consider optimizing all metrics chosen by stakeholders. While individual and group fairness might not inherently conflict as they reflect different aspects of consistent moral and political concerns (Binns, 2020), conflicts may arise in practice when optimizing specific metrics (Mehrabi et al., 2021). In our view, these approaches relate back to attempting to solve a socio-technical problem with more technical solutions, underscoring the necessity of group negotiations to resolve differing views.
6.3.3. Team Negotiation Methods for Metric Consensus
@@
In our framework, we emphasized collective metric consensus to foster mutual acceptance, and our findings demonstrate that stakeholders can reach a consensus without the interference and aid of researchers.
The varying strategies adopted for negotiating consensus within the teams of our study indicate that robust negotiating approaches need to be developed, codified, and supported by tools and processes. For example, all teams in our study used a voting mechanism, however, sometimes this led to a ’voting deadlock’. We suggest that majority-based voting or Borda voting is integrated into this process as a key stage of consensus building.
@@
In addition, it needs to be ensured that the negotiation process is fair, that minority opinions are heard, while guarding against discriminatory views. Part of the EARN framework encouraged checking mutual understandings, as they were explaining fairness metrics to each other, using the FEE tool. The negotiation process itself could be enhanced by incorporating tools such as reason-based argumentation frameworks (Jackson et al., 2023). This approach formally models moral debates, meticulously evaluates the moral acceptability of each argument, and clearly demonstrates the logical foundation and interrelations among them, thus ensuring the negotiation process is fair and inclusive.
@@ 6.3.4. Scaling Up The Process
The first part of the EARN process, the individual sessions, could be readily translated online, and used by people simultaneously without researcher involvement. This relies on the stand-alone use of the FEE tool, and choosing features and at least one fairness metric.
For team sessions, developing an online system for a larger group is essential. At the minimum, as outlined in the previous section, this needs to support some voting, as well as negotiation around fairness metrics. However, the negotiation support may quickly become unwieldy with too many participants. To overcome this, stakeholders could be chosen who represent stakeholders’ viewpoints for negotiations. This approach however places demands on how to appropriately select the right representatives.
@@ 7. Conclusions
In this paper, we proposed the EARN Fairness framework to uncover fairness metric consensus among stakeholders without prior AI knowledge. We applied our framework in a credit rating scenario with 18 participants to identify decision subjects’ personal metric preferences with their fairness threshold settings (i.e., the acceptable level of unfairness) and then, by studying three teams of 6 participants, to reveal their team consensus on metric selection. We found that:
  • For personal preferences, the top 3 metrics most favored by individuals are Conditional Statistical Parity (for subgroup fairness), Consistency, and Counterfactual Fairness.
  • Stakeholders have varying tolerance levels for unfairness and they set the most stringent standards for individual fairness.
  • For team consensus, teams typically choose to compromise by combining multiple metrics to address differences in personal preferences. Three teams uniformly show a preference for Conditional Statistical Parity at the subgroup level, yet their choices in individual fairness metrics vary.
We discussed the implications of choosing fairness metrics for credit and loan scenarios, and how to extend the EARN framework to other settings. Overall, our research presents a feasible approach for engaging stakeholders in the collective selection of AI fairness metrics.
@@ A.3. Codebooks
A.3.1. Participants’ Reasons for Selection and Exclusion
@@
A.3.2. Reasons for Fairness Threshold Settings
We analyze the reasons behind participants’ threshold settings.
Table 7. Codebook-Reasons for Fairness Threshold Settings
Code Code Description Example Pts. Freq.
@@ A.3.3. Participants’ Feedback on UI Design
User Experience Cognitively Challenging Expressing a user experience that requires a high level of mental effort and concentration. II got a little bit lost at first so until I finish the whole process. […] It makes a lot of time to understand. Get to half an hour for me. (P8) 8 12

### v2

@@ 6. Discussion
In this section, we will first discuss the limitations of our work and future research. Then, we will discuss implications for the design of fairness tools and processes by considering how UIs could be extended to express personal preferences, the complexity of accommodating diverse personal fairness preferences, how to support negotiation in groups, and how to alleviate problems in reaching consensus. We reflect on how these tools and processes would need to scale up to take into account larger and distributed stakeholder groups.
@@ 6.1. Limitations and Future Work
6.2. Implications for Human-Centered Fairness
We argue that tackling AI fairness transcends the bounds of solely technical, AI expert-driven approaches, evolving into a broader, human-centered concern. Our work demonstrates the feasibility of this approach. We will discuss how our work presenting the EARN Fairness framework and the FEE interactive systems has implications for design to support stakeholders’ elicitation of personal preferences and negotiation of consensus fairness metrics.
@@ 6.2.1. Explaining Fairness Metrics
As mentioned in Section 2, most research indirectly explores metrics that resonate with users’ fairness perceptions (Srivastava et al., 2019; Nakao et al., 2022; Stumpf et al., 2024; Saxena et al., 2019). Instead, our framework shares a similar philosophy with the work of Cheng et al. (Cheng et al., 2021), using the FEE interactive system to reveal the underlying logic of metrics, and thus enabling users to decide which fairness metric aligns best with their expectations. Although our FEE tool only covered a subsection of metrics, this could be easily extended to cover more and different fairness metrics, for a range of features extending beyond protected and legitimate features. Our tool also allows researchers to extend it to different application scenarios by changing datasets or incorporating other AI models’ fairness outcomes.
@@
6.2.2. Complexity of Diverse Fairness Preferences
@@
From the perspective of personal preferences, we observed differences between the fairness metrics preferred by stakeholders without AI expertise and those commonly used by AI experts in industry and academia, such as demographic parity, equal opportunity, and equalized odds (Xu, 2022; Bajracharya et al., 2023). Similarly, Nakao et al. (Nakao et al., 2023) found that loan officers’ preferences diverged from data scientists’. This highlights the need to involve a broader base of stakeholders in fairness metric decisions. However, even within the same type of stakeholders, in this case, decision subjects, we observed different understandings of fairness, not only in metrics but also in fairness categories. This finding echoes previous research (Lee et al., 2020; Landers and Behrend, 2023; Cheng et al., 2021).
This leads to the question of what to do with different personal preferences once we have elicited them, in order to accommodate differences. Given that there is no one-size-fits-all metric and selecting only one metric is insufficient (Mehrabi et al., 2021), one possible approach for AI practitioners and researchers, as revealed through our findings, is preference integration. When teams negotiated their preferred fairness metric, we observed a common pattern: all teams tended to merge diverse preferences to form a more comprehensive and inclusive definition of AI fairness to resolve differences, despite using different negotiation strategies. That is, in the credit rating scenario, all teams agreed on a dual focus on the absence of prejudice or favoritism toward both individuals and groups. While individual and group fairness might not inherently conflict as they reflect different aspects of consistent moral and political concerns (Binns, 2020), conflicts may arise in practice when optimizing specific metrics (Mehrabi et al., 2021). We will next discuss how negotiations around conflicting personal preferences could be supported.
6.2.3. Supporting Team Negotiation
@@
In our framework, we emphasized collective metric consensus to foster mutual acceptance, and our findings demonstrate that stakeholders can reach a consensus without the interference and aid of researchers. We observed that varying strategies were adopted for negotiating consensus within the teams of our study, indicating that robust negotiating approaches need to be developed, codified, and supported by tools and processes.
Our current interactive system, FEE, allows each user to access fairness metrics, track assessment results, and further helps them rank metrics and determine fairness thresholds. While it can be used during the group negotiation phase to review metrics together, it does not support any structured group negotiation.
We propose a series of design improvements to support better group negotiation and communication. First, tools such as FEE will need to be expanded to display personal fairness metric preferences of individuals, as shown by other tools used in algorithmic fairness (Lee et al., 2019a). This will enable stakeholders to see others’ choices, helping them better understand other positions. It remains to be investigated what the implications are of making these choices directly attributable to individuals or only making anonymous contributions visible.
Differences in choices could be made visible through visualization tools to facilitate collective decision-making. For example, tools such as graph visualization can display the proximity of individual preferences (Alonso et al., 2007), while visual ratings of support or opposition to specific metrics (Zilouchian Moghaddam et al., 2015; Kriplean et al., 2012) can quickly pinpoint areas of potential consensus or disagreement. Meanwhile, it needs to be ensured that while guarding against discriminatory views, minority opinions are heard and the negotiation process is fair.
@@
Therefore, users should be able to provide a brief justification for their selected fairness metrics. This would foster an evidence-based negotiation environment where stakeholders can engage in meaningful discussions. The negotiation process itself could then be enhanced by incorporating tools such as reason-based argumentation frameworks (Jackson et al., 2023). This approach formally models moral debates, meticulously evaluates the moral acceptability of each argument, and clearly demonstrates the logical foundation and interrelations among them, thus ensuring the negotiation process is fair, inclusive, and conducive to building consensus.
6.2.4. Handling Deadlocks and Failures during Negotiations
We also observed instances of temporary ”voting deadlock” during group negotiations but there might also be more permanent failures to reach consensus employing the EARN Fairness framework.
First, rules for building consensus could be embedded into tools that support negotiation, such as voting mechanisms (Lee et al., 2019a, b), including majority-based voting observed in our group sessions, or the Borda voting method, as implemented in the ”WeBuildAI” framework (Lee et al., 2019b). Additionally, tools could support game-rule-driven negotiations to reduce the risk of failure (Engelbutzeder et al., 2023). Gamified elements can enhance team cohesion, promote group identity, stimulate group dynamics, and improve enjoyment and user experience (Prost et al., 2019). However, when designing the ”game rules” for negotiation, it would be essential to encourage cooperative rather than competitive engagement, to avoid personal preferences becoming entrenched, akin to what we saw in Team 2 (Section 5.2).
To overcome permanent failures, we might also consider optimizing across all metrics chosen by stakeholders. To do this we can apply a preference weighting mechanism to calculate the score for each metric based on individual preferences, which automatically recommends the ”winning” metric. For example, in our case, weights of 3, 2, and 1 can be assigned for the Top-1, Top-2, and Top-3 choices, and a final weighted ranking can be displayed to the users. In our credit rating scenario, Conditional Statistical Parity (36 points) far outweighed Consistency (23 points), Counterfactual Fairness (21 points), and other metrics. However, in our view, these approaches relate back to attempting to solve a socio-technical problem with more technical solutions.
@@ 6.2.5. Scaling Up The Process
The first part of the EARN Fairness process, the individual sessions, could be readily translated online, and used by people without researcher involvement in an asynchronous and distributed manner. This relies on the stand-alone use of the FEE tool, and choosing features and at least one fairness metric.
To support the negotiation among distributed groups of people, developing an online system for a larger group is essential. This could be achieved by integrating synchronous and asynchronous online collaboration tools. For example, other tools (Zhang et al., 2023; Kawakami et al., 2024) allow personal preference data to be imported into online whiteboards to display the results to a group of stakeholders and attach reasoning or justifications, thus providing a foundation for deliberation. However, these tools will not be adequate for larger groups as negotiation support may quickly become unwieldy with too many participants. How to address this challenge is an open question but could be solved with other democratic processes already in place, for example, by choosing stakeholder representatives.
@@ 7. Conclusions
In this paper, we proposed the EARN Fairness framework to uncover fairness metric consensus among stakeholders without prior AI knowledge. We applied our framework in a credit rating scenario with 18 participants to identify decision subjects’ personal metric preferences, and then, by grouping them into three teams of 6 participants each, we studied how they negotiated consensus on metric selection. We found that:
  • Lay stakeholders differ from popular metrics in industry and academia, and there is no overall preference for one metric. There is a need for consensus to be negotiated.
  • There are many, perfectly good reasons for choosing fairness metrics which are used in negotiating fairness in groups.
  • We observed some problems in reaching consensus during the negotiations, mainly centering on ways to resolve conflicts. For consensus, groups typically chose to compromise by combining multiple metrics to address differences in personal preferences.
We discussed the implications for future tools and processes that support reaching consensus on choosing fairness metrics. Our framework and interactive system present a feasible step for engaging stakeholders in the expression of personal preferences and collective selection of metrics, fostering more equitable and inclusive AI fairness.
@@ A.3. Codebooks
A.3.1. Reasons for Fairness Threshold Settings
We analyze the reasons behind participants’ threshold settings.
Table 5. Codebook-Reasons for Fairness Threshold Settings
Code Code Description Example Pts. Freq.
A.3.2. Participants’ Reasons for Selection and Exclusion
@@ A.3.3. Participants’ Feedback on UI Design
User Experience Cognitively Challenging Expressing a user experience that requires a high level of mental effort and concentration. I got a little bit lost at first so until I finish the whole process. […] It makes a lot of time to understand. Get to half an hour for me. (P8) 8 12